OpenStudy (redneckoutlaw):

What is the exact value of 14 times z squared all over the square root of z to the ninth power.?

OpenStudy (imstuck):

\[\frac{ 14z ^{2} }{ \sqrt{z ^{9}} }\]is that it?

OpenStudy (redneckoutlaw):

yes

OpenStudy (redneckoutlaw):

whats the exponiet above the bottom Z?

OpenStudy (imstuck):

\[\frac{ 14z ^{2} }{ z ^{\frac{ 2 }{ 9 }} }\]Do you know the rules about dividing exponents?

OpenStudy (imstuck):

Here is how you find the exponent from a square root:

OpenStudy (redneckoutlaw):



OpenStudy (imstuck):